# About testing Bell locality at colliders

### v2

@@ Abstract.
Colliders enable the testing of quantum mechanics at its most fundamental level, and in the presence of both strong and electroweak interactions. Particle physics naturally provides examples of systems that consist not only of qubits but also qutrits. Quantum tomography at colliders makes possible the measurement of entanglement and the testing of Bell locality. We provide a comprehensive explanation of how this is achieved and why it offers a novel perspective on the loopholes exploited by local hidden variable models to evade Bell tests.
@@ 1. Introduction
One of the most distinctive features of a quantum system is the possibility of entanglement, which refers to quantum correlations among observables that defy classical physics explanations [1, 2, 3, 4]. Entanglement often clashes with common sense, which is rooted in local realism. Consequently, classical models encompassing standard quantum mechanics have been proposed to explain phenomena in terms of underlying hidden variables. These variables can be deterministic or stochastic, but to replicate the probabilistic nature of quantum mechanics, only their average over a specific distribution influences the outcome of measurements in both cases.
@@
J. Bell’s seminal paper [28] provided a new perspective to the discussion. It demonstrated that a significant portion of these contextual models of hidden variables, those adhering to a locality condition (Bell locality), satisfy specific constraints (Bell inequalities) that can be violated by quantum mechanics [29, 30]. This result enables the differentiation between classical models and quantum mechanics through actual experiments [31, 32, 33, 34, 35]. Numerous tests of this nature have been conducted using diverse systems, spanning low energies within atomic and solid-state physics [36, 37, 38, 39, 40, 41, 42, 43] and, more recently, at high energies in collider experiments (for a comprehensive review, refer to [44]).
Colliders are the only settings where quantum mechanics is tested at its most fundamental level, in the presence of both strong and electroweak interactions. Particle physics naturally provides examples of systems that consist not only of qubits but also qutrits.
At particle colliders, it is possible to reconstruct the properties of the particles that produced in a collision. The spin state of the products of a scattering process can be determined by analyzing the angular distributions of the momenta of the related final-state particles, as recorded by a detector. This process, known as quantum state tomography, provides a complete description of the state. It is important to note that this approach differs from the traditional Bell inequality test, where only joint probabilities along specific directions in space are measured. For the latter, since only these probabilities are measured, the final state is only partially determined and this lack of knowledge can be exploited to devise loopholes in the test. At colliders, the complete knowledge of the state appears to close all loopholes, making the testing of Bell inequalities a straightforward determination of suitable observables on the reconstructed state. To avoid confusion about the meaning of this procedure, we thoroughly examine the testing of nonlocality at colliders and compare it to conventional low-energy experiments.
We are aware that the ideal configuration for testing nonlocality against hidden variable models would be a black box in which classical inputs can be varied at will and classical outputs recorded [45, 46]. This approach has been extensively used in quantum information theory [47, 48, 30] (in particular in the context of quantum key distribution protocols [49, 50]) to avoid a precise description of the employed equipment, which is often unavailable or introduces uncertainties due to imperfect measurement settings [51]. The S-matrix approach to quantum field theory was originally inspired by the same principle of only using the inputs (incoming particles) and outputs (outgoing particles) in those cases, as for the strong interactions, for which there was only limited understanding. Unfortunately, no present or planned collider experiment satisfies the conditions for device independence. Consequently, we confine our discussion to nonlocality within a device-specific context. In the language of quantum information theory, the measurements used in testing Bell inequalities, as those discussed in this work, must be understood as nonlocality quantum witnesses rather than proper Bell tests [30].
@@
    This is followed by a discussion of how non-commuting variables are reconstructed from commuting ones;
    We then provide an overview of quantum tomography and explore to what extent it relies on quantum mechanics;
    Finally, we explore how the discussion of loopholes in Bell tests is reshaped within the context of quantum tomography.
The paper concludes with a concise overview of the current research on Bell nonlocality at colliders.
@@ 2. Quantum entanglement
By taking spin measurements along different directions, the state is not defined by the conservation of the angular momentum and one detects correlations not amenable to a classical explanation. In fact, perfect knowledge of the joint spin state of the particles, given in (2.1), results in a total indetermination of the state of each of the constituent particles. Since this would occur even if the two particles were separated at an arbitrarily large distance at the time of the measurement, one speaks of quantum nonlocality.
@@ 3. Bell nonlocality
where $\sigma_{i}$ are the Pauli matrices, $\bf 1$ is the unit $2\times 2$ matrix, and where the sums involving the indices $i$ and $j$ run over the labels $x$, $y$ and $z$, representing any orthonormal spatial reference frame. The real coefficients $B_{i}^{(+)}={\rm Tr}[\rho\,(\sigma_{i}\otimes{\bf 1})]$ and $B_{j}^{(-)}={\rm Tr}[\rho\,({\bf 1}\otimes\sigma_{j})]$ give the spin averages (i.e. the polarizations) of the corresponding particles, while the real matrix $C_{ij}={\rm Tr}[\rho\,(\sigma_{i}\otimes\sigma_{j})]$ contains the spin correlations. Note that while the density matrix in (3.5) is normalized to ${\rm Tr}[\rho]=1$, extra constraints on $B_{i}^{(+)}$, $B_{j}^{(-)}$ and $C_{ij}$ are needed to guarantee non-negative eigenvalues. These extra conditions are in general non-trivial, requiring all principal minors of $\rho$ to be non-negative. Knowledge of $\rho$ allows to compute the average of any two-spin observable $\mathcal{O}$ as $\langle\mathcal{O}\rangle={\rm Tr}[\rho\,\mathcal{O}]$, including the probabilities appearing in (3.4).
@@
In order to put under test this Bell inequality, one needs to experimentally determine the matrix $C$ and then suitably choose four spatial directions ${\vec{n}}_{1}$, ${\vec{n}}_{2}$, ${\vec{n}}_{3}$ and ${\vec{n}}_{4}$ that maximize the left-hand side of Eq. (3.9). Fortunately, this maximization process can be easily performed analytically, for a generic spin correlation matrix. This has been dimonstrated in [54]. Indeed, consider the matrix $C$ and its transpose $C^{T}$ and form the symmetric, positive, $3\times 3$ matrix $M=C^{T}C$; its three eigenvalues $m_{1}$, $m_{2}$, $m_{3}$ can be ordered in decreasing order: $m_{1}\geq m_{2}\geq m_{3}$. Then, the following result holds: the two-spin state $\rho$ in Eq. (3.5) violates the inequality (3.9), or equivalently (3.4), if and only if the sum of the two largest eigenvalues of $M$ is strictly larger than 1, that is
@@
It is important to note that the above formulation diverges from the conventional approaches commonly employed in testing Bell inequalities, particularly within the field of quantum optics. While in standard tests, the probabilities that enter the inequality (3.4) are experimentally determined by measuring the frequencies of events for each choice of the directions ${\vec{n}}_{i}$, in the above approach, actual measurements of probabilities are avoided. Instead, we rely on the experimental determination of the spin correlation matrix $C$. The entries of this matrix are typically reconstructed using suitable angular distributions observed at collider experiments, without any specific reference to predefined sets of directions ${\vec{n}}_{i}$.
@@ 4. Quantum information observables in the $S$-matrix framework
A superficial comparison with the low-energy tests of quantum mechanics based on polarized photons and polarimeters might suggest that, in order to explore quantum entanglement and Bell inequality violation at colliders, one must necessarily use polarized beams and polarized targets. Yet this is not the way high-energy physics has developed and there are essentially no experiments using such beams and targets.
@@
Almost all experiments are, instead, based on (mostly) unpolarized beams and detectors that do not measure directly polarizations. Is this situation a show stopper [^1] See, for example, [55]. for testing quantum mechanics—be it Bell nonlocality or entanglement—at colliders? Not at all—as the rich physics program on polarization effects at colliders shows. Polarizations are reconstructed from the angular distributions of particles in the final state. As explained in detail below, the momentum angular distribution is used to estimate correlations, thereby bringing us from the cross section to a density matrix. It is in the space in which this matrix is defined that entanglement and Bell inequality violation must be studied—not in the cross section. This is a powerful way of studying polarizations and it allows for their measurement also of those particles that are not directly seen in the detector because they decay before reaching it.
@@
in which $\expectationvalue{s_{i}}$ are the polarization projections onto a given basis.
@@
with $\theta_{i}$ the angles described by the momenta of the charged pions in the decay of the $\tau$ lepton. This reconstruction, in which the decaying particle acts as its own polarimeter, is based on the independent experimental determination of the polarimetric vector, the coefficient $\alpha_{i}$ in Eq. (4.5), which turns out to be $\alpha_{i}=1$ for the single-pion $\tau$ decay. This determination is mirrored in low-energy experiments with photons by the measurement of the properties of the polarimeter used in the experimental setup. Both experiment types require an external input which however comes from experiments that are independent of those testing the Bell inequality. Equation (4.5), apart for the polarimetric vector, is nothing but the result of the conservation of angular momentum in the decay—no quantum mechanics is required.
Figure 2. Clauser’s experimental setup (Stuart Freedman, Berkeley [$C$ BY-NC 4.0]) and the CMS detector at the LHC (CERN [$C$ BY-NC 4.0]). Neither is as simple as a falling apple and the device-independence of potential Bell test has to be checked on a case-by-case basis.
All experiments—be them at low or high energy (see Fig. 2)—must necessarily assume some model about the working of the polarimeter utilized in the measurements. It is the polarimeter that transforms a measured data (light or darkness in the case of photons, angular distribution of momenta in the case of the decay of a particle) into a polarization measurement. In the case of the atomic-physics tests, this equipment consists of the photon polarimeter, while in the case of the high-energy tests, in the polarimetric vector and spin-analyzing power of a particle emitted in a decay. In both cases, the actual operation of the polarimeter has to be determined independently of the Bell test or a measurement of entanglement, in separate experiments providing the characteristic properties of the photon polarimeter or of the polarimetric vector. It is only after this step that the experimentally collected data—be them in the form of frequencies, probabilities or correlations—can be presented as those giving photon or elementary particle polarizations.  [^2] On the other hand, in an analytic computation or a Monte Carlo simulation, the operation of the polarimeter can be determined directly from the specific features of the model used, rather than being derived from some other data. For instance, in the context of tests at colliders, the polarimetric vector in the fully leptonic decay of the top quark is found to be equal to one due to the chiral coupling between elementary fermions and gauge bosons in the Standard Model.
@@
If the reader feels uneasy about using commuting variables, such as momenta, to reconstruct a non-commuting variable like the spin, consider that most, if not all, non-commuting variables are reconstructed using commuting variables. For instance, in the Stern-Gerlach experiment, the spin of a neutron is measured by its position (or momentum) on a screen. The commuting variable simply provides the spin value, which is either up or down. It is only through repeated experiments using different directions and comparing the results for the neutron’s spin that its non-commutative nature is revealed.
Similarly, at colliders, specific momentum angular distributions are used as distribution of probabilities to estimate correlations, thereby bringing us from the cross section to the density matrix. It is in the Hilbert space on which this matrix acts that entanglement and Bell inequality violation must be studied—not in the cross section.
@@
$|\Psi\rangle=\dfrac{1}{\sqrt{2}}\Big{(}|\tau^{-}_{R}\rangle|\tau^{+}_{L}\rangle+|\tau^{-}_{L}\rangle|\tau^{+}_{R}\rangle\Big{)}\$, (4.10)
@@
in the case of Eq. (4.9). Once the state is known, tests of Bell locality or any other property are just a matter of taking expectation values of the appropriated operators. Here we mimic those leading to the traditional setting of a Bell test, by studying the inequality (3.4).
@@
This derivation explicitly shows how the Bell inequality can be tested at a collider and how the non-separable, Bell state violates the inequality. It follows the traditional protocol of fixing some directions and comparing probabilities. With quantum tomography this is only an afterthought since the reconstruction of the state automatically shows the presence or absence of entanglement and the implied Bell test outcome—having the state we can perform any test we wish on it to highlight its quantum properties.
5. Quantum tomography: a new take on loopholes

### v3

# About witnessing Bell non-locality at colliders

@@ Abstract.
High-energy colliders enable the testing of quantum mechanics at its most fundamental level, in the presence of strong and electroweak interactions, with systems that consist of qubits (fermions) and qutrits (massive spin-1 bosons). Quantum state tomography at colliders enables the witnessing of entanglement and Bell non-locality, two defining characteristics of quantum mechanics. We offer a comprehensive explanation of the underlying principles and the methods employed to achieve this remarkable feat.
@@ 1. Introduction
One of the most unique characteristics of a quantum system is the possibility of entanglement—which refers to the presence of quantum correlations among observables that defy classical physics explanations [1, 2, 3, 4]. Entanglement often clashes with common sense, which is rooted in local realism. Consequently, classical models encompassing standard quantum mechanics have been proposed to explain phenomena in terms of underlying hidden variables. These variables can be deterministic or stochastic, but to replicate the probabilistic nature of quantum mechanics, only their average over a specific distribution influences the outcome of measurements in both the cases.
@@
J. Bell’s seminal paper [28] provided a new perspective to the discussion. It demonstrated that a significant portion of these contextual models of hidden variables, those adhering to a locality condition (Bell locality), satisfy specific constraints (Bell inequalities) that can be violated by quantum mechanics [29, 30]. This result makes it possible to distinguish between classical models and quantum mechanics through actual experiments [31, 32, 33, 34, 35]. Numerous tests of this nature have been conducted using diverse systems at low energies, for instance within atomic and solid-state physics [36, 37, 38, 39, 40, 41, 42, 43]. More recently, entanglement and Bell locality have also been investigated at the higher energies probed in collider experiments (for a comprehensive review, see [44]). This unique setting makes it possible to test quantum mechanics at its most fundamental level, in the presence of both strong and electroweak interactions, as well as with systems that consist not only of qubits, but also of qutrits.
At particle colliders, it is possible to reconstruct the properties of the particles produced in a collision. For instance, the spin state of the products of a scattering process can be determined by analyzing the angular distributions of the momenta of the related final-state particles, as recorded by a detector. This process, known as quantum state tomography, provides a complete description of the state. Consequently, witnessing Bell non-locality in a bipartite system at colliders is as simple as determining the expectation value of suitable observables on the reconstructed state.
To avoid confusion about the meaning of this procedure, it is important to bear in mind the difference between a true Bell test and a Bell non-locality witness that quantum state tomography makes available. The ideal configuration for testing the presence of non-locality is a black box in which classical inputs can be varied at will (the measurements setting), and classical outputs recorded (the measurement outcomes) [45, 46]. In this case, measuring a violation of the Bell inequality would support the presence of genuine Bell non-locality in the system. This would then hinder hidden variable models that could recover the experimental result only through possible loopholes specific to the experiment setup. The black box approach has been extensively used in quantum information theory [47, 48, 30], in particular in the context of quantum key distribution protocols [49, 50]. This approach avoids providing a precise description of the employed equipment, which is often unavailable and would introduce uncertainties due to imperfect measurement settings [51]. Though most low-energy experiments are often device dependent, the black box paradigm can be recovered after the working of their experimental settings (for example, the polarimeters and the source generating the system of interest) is certified through dedicated experiments.
The S-matrix approach to quantum field theory was originally inspired by the same black-box principle using the inputs, incoming particles, and outputs, outgoing particles. This proved fruitful in modeling processes for which the ongoing dynamics was only limitedly understood, as it was the case for the strong interactions. At colliders, spin correlations are reconstructed via a series of tomographic measurements which use, as input, the angular distribution of the momenta of suitable final state particles that reveal the spin of the progenitor particles. The outcome of this procedure is the spin state of the quantum system under investigation, and the presence of properties such as entanglement or Bell non-locality can then be probed through the density matrix formalism with dedicated observables. Accordingly, the results of a collider experiment will satisfy the conditions for device independence only if the collision dynamics, the detector performance and the reconstruction procedure of quantum correlations are all independently certified. Although this could be—and often it has already been—achieved through a series of independent experiments, the ideal setting of a black-box for performing Bell tests cannot be achieved due to the lack of tunable classical inputs that specify, for instance, the measurement settings. As in collider experiments the settings are essentially chosen by the quantum system under investigation itself, it is necessary to introduce an element of trust on the experimental platform, which makes it impossible to produce certified device-independent results. In the language of quantum information theory, the measurement of Bell non-locality produced at collider experiments via quantum tomography are to be understood as quantum witnesses of non-locality rather than the results of proper Bell tests [29, 30] achievable only in with a black-box setting.
@@
    This is followed by a discussion of how non-commuting observables are reconstructed from commuting ones;
    We then provide an overview of quantum tomography—stressing the strengths and weaknesses of the procedure;
    Finally, we comment on the actual implementation of hypothetical local hidden models within the framework of a fundamental theory.
    In the Appendix, we explore how the discussion of traditional loopholes is reshaped within the context of collider experiments and quantum tomography.
The paper final section contains a concise overview of the current possibilities offered by quantum state tomography at colliders.
@@ 2. Quantum entanglement
By taking spin measurements along different directions, the state is not defined by the conservation of the angular momentum and one detects correlations not amenable to a classical explanation. In fact, perfect knowledge of the joint spin state of the particles, given in (2.1), results in a total indetermination of the state of each of the constituent particles. Since this would occur even if the two particles were separated at an arbitrarily large distance at the time of the measurement, one speaks of quantum non-locality.
@@ 3. Bell non-locality
where $\sigma_{i}$ are the Pauli matrices, $\bf 1$ is the unit $2\times 2$ matrix, and where the sums involving the indices $i$ and $j$ run over the labels $1$, 22 and $3$, representing any orthonormal spatial reference frame. The real coefficients $B_{i}^{(+)}={\rm Tr}[\rho\,(\sigma_{i}\otimes{\bf 1})]$ and $B_{j}^{(-)}={\rm Tr}[\rho\,({\bf 1}\otimes\sigma_{j})]$ give the spin averages (i.e. the polarizations) of the corresponding particles, while the real matrix $C_{ij}={\rm Tr}[\rho\,(\sigma_{i}\otimes\sigma_{j})]$ contains the spin correlations. Note that while the density matrix in (3.5) is normalized to ${\rm Tr}[\rho]=1$, extra constraints on $B_{i}^{(+)}$, $B_{j}^{(-)}$ and $C_{ij}$ are needed to guarantee non-negative eigenvalues. These extra conditions are in general non-trivial, requiring all principal minors of $\rho$ to be non-negative. Knowledge of $\rho$ allows to compute the average of any two-spin observable $\mathcal{O}$ as $\langle\mathcal{O}\rangle={\rm Tr}[\rho\,\mathcal{O}]$, including the probabilities appearing in (3.4).
@@
In order to put under test the Clauser-Horne-Shymony-Holt (CSHC) inequality [54] in Eq. (3.9) [^1] This inequality differs from the original Bell inquality by being about correlations rather than probabilities., one needs to experimentally determine the matrix $C$ and then suitably choose four spatial directions ${\vec{n}}_{1}$, ${\vec{n}}_{2}$, ${\vec{n}}_{3}$ and ${\vec{n}}_{4}$ that maximize the left-hand side of Eq. (3.9). Fortunately, this maximization process can be easily performed analytically, for a generic spin correlation matrix. This result has been demonstrated in [55]. Indeed, consider the matrix $C$ and its transpose $C^{T}$ and form the symmetric, positive, $3\times 3$ matrix $M=C^{T}C$; its three eigenvalues $m_{1}$, $m_{2}$, $m_{3}$ can be ordered in decreasing order: $m_{1}\geq m_{2}\geq m_{3}$. Then, the following result holds: the two-spin state $\rho$ in Eq. (3.5) violates the inequality (3.9), or equivalently (3.4), if and only if the sum of the two largest eigenvalues of $M$ is strictly larger than 1, that is
@@
As mentioned in the introduction, it is important to stress how the above procedure differs from the conventional approach commonly employed in testing Bell inequalities, particularly within the field of quantum optics. In standard Bell tests, the probabilities that enter the inequality (3.4) are experimentally determined by measuring the frequencies of events for each choice of the directions ${\vec{n}}_{i}$. In the above approach, instead, we only rely on the experimental determination of the spin correlation matrix $C$ provided by quantum state tomography. As no experimental setting can be arbitrarily varied, it is necessary to trust that the angular distributions used to reconstruct the correlation matrix are produced by a fair source. To the extent that in quantum tomography both outcomes and the settings are determined together and not independently from each other, this approach provides, in the language of the quantum information theory, a witness of Bell non-locality and not a Bell test.
@@ 4. Quantum information observables in the $S$-matrix framework
A superficial comparison with the low-energy tests of quantum mechanics based on polarized photons and polarimeters might suggest that, in order to explore quantum entanglement and Bell non-locality violation at colliders, one must necessarily use polarized beams and polarized targets. Yet this is not the way high-energy physics has developed and there are essentially no experiments using such beams and targets.
@@
Almost all experiments are, instead, based on (mostly) unpolarized beams and detectors that do not measure directly polarizations. Is this situation a show stopper [^2] See, for example, [56]. The argument in this paper hinges on the existence of a local hidden variable model—introduced by Kasday[57]—in which the hidden variables are directly related to the momenta and their distribution given by the Standard Model differential cross section. As explained in this work, this implies that the model is an instance of the class of super-deterministic models which are discussed in Section 5. for testing quantum mechanics—be it Bell non-locality or entanglement—at colliders? Not at all—as the rich physics program on polarization effects at colliders shows. Polarizations are reconstructed from the angular distributions of particles in the final state. As explained in detail below, the momentum angular distribution is used to estimate correlations, thereby bringing us from the cross section to a density matrix. It is in the space in which this matrix is defined that entanglement and Bell non-locality must be studied—not in the cross section. This is a powerful way of studying polarizations and it allows to measure also the polarization of particles that are not directly seen in the detector because they decay sufficiently fast. Studies of entanglement and Bell non-locality in isospin space, as reported in [58], further support the notion that quantum mechanics can indeed be tested using quantum state tomography. Notably, this approach eliminates the need for momentum measurements, and the final state is entirely determined by the values of the branching rates in the various channels.
@@
in which $\expectationvalue{s_{i}}$ are the average polarization components in a given basis.
@@
with $\theta_{i}$ the angles described by the momenta of the charged pions in the decay of the $\tau$ lepton.
This reconstruction, where the decaying particle acts as its own polarimeter, relies on the independent experimental determination of the polarimetric vector. It is commonly performed for hyperons and other heavy baryons in charmonium decays [59, 60, 61]. This reconstruction may not be available from the data alone for particles that are not produced in a definite state of polarization. This is the case of the $\tau$ particle at present colliders, for which the coefficient $\alpha_{i}$ in Eq. (4.5) turns out to be $\alpha_{i}=1$ for the single-pion $\tau$ decay, solely due to the conservation of angular momentum in the decay process.
Figure 2. Clauser’s experimental setup (Stuart Freedman, Berkeley [$C$ BY-NC 4.0]) and the CMS detector at the LHC (CERN [$C$ BY-NC 4.0]). Neither is as simple as a falling apple and the device-independence of potential Bell test has to be certified.
All experiments—be them at low or high energy (see Fig. 2)—must necessarily rely on working of the polarimeter utilized in the measurements. It is the polarimeter that transforms a measured data (light or darkness in the case of photons, angular distribution of momenta in the case of the decay of a particle) into a polarization measurement. In the case of the atomic-physics tests, this equipment consists of the photon polarimeter, while in the case of the high-energy tests, in the polarimetric vector and spin-analyzing power of a particle emitted in a decay. In both cases, the actual operation of the polarimeter has to be certified in independent experiments apt to provide the characteristic properties of the photon polarimeter or of the polarimetric vector. [^3] It is worth pointing out that in an analytic computation, or a Monte Carlo simulation, the operation of the polarimeter can be determined directly from the specific features of the model used, rather than being derived from some other data. For instance, in the context of tests at colliders, the polarimetric vector in the fully leptonic decay of the top quark is found to be equal to one due to the chiral coupling between elementary fermions and gauge bosons in the Standard Model. It is only after this step that the experimentally collected data—be them in the form of frequencies, probabilities or correlations—can be processed to create device-independent results. In regard of this, experiments always compare models against the data. The data represent the “real world,” and when presented with a model, we assess how well the data aligns with it. However, there is no direct way to test the “real world.” Consequently, the Bell test simply compares the factorization of probabilities (Bell locality) against the data. Its results can be used to rule out local theories (if the inequality is violated) and even quantum mechanics (if the violation exceeds a critical value known as Tsirelson’s bound [62]). The same principle applies to measurements of entanglement.
@@
If the reader feels uneasy about using commuting observables, such as momenta, to reconstruct a non-commuting observable like the spin, consider that observed quantities are always real numbers and, as such, commuting. For instance, in the Stern-Gerlach experiment, the spin of a neutron is measured by its position (or momentum) on a screen. The commuting variable simply provides the spin value, which is either up or down. It is only through repeated experiments using different directions and comparing the results for the neutron’s spin that its non-commutative nature is revealed.
Similarly, at colliders, specific momentum angular distributions are used as distribution of probabilities to estimate correlations, thereby bringing us from the cross section to the density matrix. It is in the Hilbert space on which this matrix acts that entanglement and Bell non-locality must be studied—not in the cross section.
@@
[…]
@@
in the case of Eq. (4.9). [^5] This is an essential point: the helicities (or the polarizations), which are real numbers and therefore commuting quantities, only show their non-commuting properties after being organized in the corresponding Hilbert space in the states in Eqs. (4.13)–(4.14). Even if the cross section were to explicitly depend on the spin variable and the detector contained a polarized target, measurements could still rely on the final momentum distribution to reconstruct the direct effect of the spin interactions. Once the state is known, tests of Bell locality or any other property are just a matter of taking expectation values of the appropriated operators. Here we mimic those leading to the traditional setting of a Bell test, by studying the inequality (3.4).
@@
This derivation explicitly shows how Bell non-locality can be probed at a collider experiment and how the non-separable, Bell state violates the inequality. It follows the traditional protocol of fixing some directions and comparing probabilities. With quantum tomography this is only an afterthought since the reconstruction of the state automatically shows the presence or absence of entanglement and the implied Bell non-locality—having the state we can perform any test we wish on it to highlight its quantum properties.
5. Quantum state tomography and hidden variables in collider experiments
The investigation of non-locality performed with quantum state tomography cannot be regarded as a Bell test as it is inherently open to a fundamental loophole. This arises from the lack of control on the angular directions used by the tomographic procedure and consists in assuming the existence of a hidden variable distribution that contains the outcome of the experiment that is about to be performed. Alternatively, this distribution could be set to match the Standard Model predictions [57], which are known to agree with experiments very well at the current precision. In either case, for particle physics experiments, this would imply that the hidden variables regulate the angular distribution of the momenta used by the tomographic procedure, in such a way that local hidden variable models could mimic the effects of a genuine non-local behavior. Furthermore, the same variables must also convey information about the detector, such as the direction of the magnetic field used to differentiate the charges of particles, due to the contextuality requirement [72].
As the collision dynamics has been certified to follow the principles of quantum field theory, all the information about the final momentum distribution and the experimental setup used must be encoded in the initial state formed by the colliding particles. This loophole then represents an instance of super-determinism, a philosophical stance that suggests in this case that hidden variables determine both the behavior (encoded in probability distributions) of the particles as well as the choices made by experimenters in a traditional Bell test. All Bell tests are inherently susceptible to this loophole [73] because the freedom in choosing the measurement directions—the free will of experimenters, in other words—is one of the assumptions underlying the derivation of the Bell inequality.
Super-deterministic models, even in their milder version in which the free will is not involved and only probabilistic distributions are to be determined beforehand, cannot be ruled out with scientific methods, neither at collider experiments nor with more traditional setups using atoms or solid-state devices. Yet, they appear to be incompatible with our understanding of the physical world as they necessitate the absolute determination of the initial conditions [74] which, in turn, requires specifying all the infinitely many digits comprising the real numbers that describe these conditions [75]. That a complete determination of the latter is necessary follows from the results of [76, 77], in which it is argued that even in the presence of strong dependence of the measurement choices on hidden variables—as it is the case in the setup of collider experiments–only an infinitesimal amount of randomness is necessary for genuine quantum non-locality to result in a violation of the Bell inequality. The initial conditions must then be fully specified for hidden variable models to overshadow the quantum effects and the vast amount of information required is to be confined within a finite volume of spacetime, thereby violating the Bekenstein bound [78]. This fundamental principle, stating that initial conditions cannot be specified with arbitrary precision, is often overlooked in discussions about determinism. [^6] The burden for those who want to revive super-determinism would be to demonstrate that the number of digits required to set the initial conditions and enable predictability for a sufficiently long period (e.g., the age of the Universe), thereby also removing the however small amount of randomness required to violate appropriate Bell inequalities, is sufficiently small to fall below the Bekenstein bound. We also remark that even if one were to disregard the above argument and insist on entertaining such a philosophical stance, the contextuality requirement would still render full-blown super-determinism models rather unappealing.
6. If they exist, hidden variables will be new physics
Hidden variables were initially proposed to complement theories that were not fundamental, such as atomic physics. On the other hand, particle physics delves into the most fundamental level of nature, where there is no room for hiding. This leads to speculations about whether the incompleteness of quantum mechanics can be rephrased as the incompleteness of the Standard Model interactions and particle content. While it is possible that a clever hidden variable model could reproduce the states identified by the tomographic procedure, regardless of the model, it will be contextual and non-local unless we consider the super-determinism choice, which we reject for the required infinite amount of information. If we attempt to construct a fundamental hidden variable model (an example of which, though non-relativistic, is Bohm’s model [79]), we will face the challenge of doing so at the microscopic level and, therefore, by means of new particles or interactions that carry the information encoded in the hidden variables. These features cannot remain hidden in a fundamental theory and must manifest themselves as these new (currently elusive) particles or interactions with special properties. From the perspective of particle physics, the search for hidden variable models can then be framed as the search for physics beyond the Standard Model.
As we contemplate the future of quantum tomography at colliders, two unexpected outcomes can emerge. If the experimental data indicated correlations surpassing the Tsirelson bound [62], the result would imply that non-locality surpasses the maximal amount that quantum mechanics can produce. All the tests conducted at colliders have so far confirmed quantum mechanics. Alternatively, the reconstructed state might not be an acceptable state within the Hilbert space of quantum mechanics. This could occur, for instance, if the experimental values of certain cross sections violate unitarity. In such cases, the problem would be addressed by seeking the new physics effects responsible for the violation of unitarity rather than being taken as a signal of the failure of quantum mechanics.
7. Outlook
In this paper, we have introduced a Bell non-locality study tailored to the spin correlations studied in collider experiments by means of quantum state tomography. The method allows for the reconstruction of the quantum state of the system under scrutiny and provides a witness of Bell non-locality. The collider results pertaining to Bell non-locality are inherently device-dependent owing to the experimental platform, which is characterized by a vulnerability reminiscent of the super-determinism loophole. As for the latter, we have stressed that super-determinism requires an absolute determination of the initial conditions incompatible with known fundamental limits in physics.
Quantum state tomography offers new insights into the fundamental workings of quantum mechanics. Its methods have been successfully applied to analyze experimental data from colliders. Specifically, analyses of $B$-meson decays at Belle II and LHCb [80], Charmonium decays at BESIII [81, 61] and pions from kaon decays [58] have been reinterpreted to achieve this. These processes have demonstrated the possibility of witnessing Bell non-locality with a significance exceeding $5\sigma$ thus extending the low-energy findings to encompass weak and strong interactions, as well as qutrits systems. Furthermore, the detection of entanglement in top-quark pair production data at the LHC has been achieved with a significance of $5\sigma$ [82, 83, 84].
Analyses using $\tau$ lepton pairs [85, 86], top-quark pairs [87, 88, 89, 90, 91, 92, 93] and vector gauge bosons [94, 95, 96, 97, 98] show that similar results can be expected in the upcoming years for many other processes. The prospects of studying Bell non-locality at future colliders has also been discussed [99, 97, 100, 98, 101, 102].
Appendix A Quantum tomography: a new take on loopholes